PeasOfClique:

sammy eats 3 apples each day, how many apple would he has eaten if he eats for 30 more days with 20 already eaten.

PeasOfClique:

*how many apples would he eat in total if he already ate 20 apple and plans to eat for 30 more days?

snowDoggie:

I cant do math

maths89898:

what are the options?

snowDoggie:

lionel

snowDoggie:

i cant do math

miazzzzlife44:

i dont f ing know

maths89898:

my guess is 99

snowDoggie:

i go with the smart lady

maths89898:

If 20 apples were already eaten that would be 3 x 6 = 18 which is 3 days so 30 - 3 = 27 x 3 = 81 + 18 = 99

maths89898:

i could be wrong tho cuz idk the options

miazzzzlife44:

me too

maths89898:

and then add two maybe so 101

miazzzzlife44:

u smart